OpenStudy (anonymous):

If f(x) = 2x−5, then what is the value of f(2) + f(5) ?

OpenStudy (compassionate):

f(2) = 2(2) - 5 f(5) = 2(5) - 5 Now, just solve. 4 - 5 10 - 5 1 5 1 + 5 = ?

OpenStudy (cggurumanjunath):

4-5=-1 not +1@Compassionate

OpenStudy (anonymous):

thank u cggurmanjunath cause i was lookin at it like that cant be rite

OpenStudy (cggurumanjunath):

YOU'RE WELCOME @nathaniellanda17 SO WHAT IS THE FINAL ANSWER ?

OpenStudy (anonymous):

4

OpenStudy (cggurumanjunath):

U R RIGHT :)